### v1

# Unifying Appearance Codes and Bilateral Grids for Driving Scene Gaussian Splatting

Abstract
Neural rendering techniques, including NeRF and Gaussian Splatting (GS), rely on photometric consistency to produce high-quality reconstructions. However, in real-world scenarios, it is challenging to guarantee perfect photometric consistency in acquired images. Appearance codes have been widely used to address this issue, but their modeling capability is limited, as a single code is applied to the entire image. Recently, the bilateral grid was introduced to perform pixel-wise color mapping, but it is difficult to optimize and constrain effectively. In this paper, we propose a novel multi-scale bilateral grid that unifies appearance codes and bilateral grids. We demonstrate that this approach significantly improves geometric accuracy in dynamic, decoupled autonomous driving scene reconstruction, outperforming both appearance codes and bilateral grids. This is crucial for autonomous driving, where accurate geometry is important for obstacle avoidance and control. Our method shows strong results across four datasets: Waymo, NuScenes, Argoverse, and PandaSet. We further demonstrate that the improvement in geometry is driven by the multi-scale bilateral grid, which effectively reduces floaters caused by photometric inconsistency. Our code is open-sourced at https://bigcileng.github.io/bilateral-driving.
Figure 1: (a) Appearance codes rely on global affine transformations but have limited modeling capability. (b) Bilateral grids perform pixel-wise transformations, improving photometric consistency but are challenging to optimize. (c) The proposed multi-scale bilateral grid unifies appearance codes and bilateral grids, enabling patch-wise transformations.
1 Introduction
Neural rendering techniques, such as NeRFs  mildenhall2021nerf ; wang2021neus ; yu2022monosdf ; liu2024rip ; yuan2024slimmerf and Gaussian Splatting (GS) kerbl20233d ; chen2024pgsr ; cheng2024gaussianpro ; zhang2024drone ; song2024sa , have demonstrated significant potential in producing high-quality 3D reconstructions by leveraging photometric consistency. However, ensuring photometric consistency across images in real-world scenarios remains a challenge, as lighting conditions, viewpoints, and camera settings can vary considerably martin2021nerf ; zhang2024gaussian ; kulhanek2024wildgaussians ; wang2025semantic . In autonomous driving applications yan2024renderworld ; wu2023mars ; tonderski2024neurad ; chen2024omnire ; li2024uniscene ; jin2024tod3cap ; tian2023unsupervised , these challenges are amplified due to the presence of multiple cameras capturing the same scene from different angles over time. As illustrated in the left half of Teaser. 1, we observe that these variations—whether temporal (across different time steps) or inter-camera (across different viewpoints)—can introduce significant discrepancies in image appearance. These inconsistencies pose difficulties in accurately reconstructing dynamic driving environments.
To address this issue, appearance codes tonderski2024neurad ; martin2021nerf ; chen2024omnire have been introduced to encode per-image information and assist in correcting the photometric discrepancies. These codes, while effective, have limitations in modeling capability since they apply a single transformation across the entire image, often neglecting local variations that may occur in complex scenes. For example, in dynamic scenes, such as those encountered in autonomous driving environments, lighting and viewpoints can change across different camera angles and time frames. As shown in the right top of Teaser. 1, a global affine transformation based on a single appearance code may not be sufficient to handle such variations (In Teaser. 1 NuScenes (a), the fence is blurred).
Recent advancements wang2024bilateral ; gharbi2017deep have introduced bilateral grids to perform pixel-wise transformations, enabling improved photometric consistency by allowing more localized adjustments. However, these grids face significant challenges in optimization, as they require complex constraints to avoid instability during training (As shown in Teaser. 1 (b), bilateral grids fail to effectively optimize the reconstruction of large, complex scenes.). In this work, we propose a novel solution—a multi-scale bilateral grid—integrating the strengths of appearance codes and bilateral grids. As shown in Teaser. 1, this new approach allows for patch-wise transformations. Interestingly, in extreme cases, the multi-scale bilateral grid naturally converges to either the bilateral grid or appearance code, depending on the scale. At the finest scale, the multi-scale grid behaves like a traditional bilateral grid, performing pixel-wise transformations to adjust for local variations. In contrast, at the coarsest scale, it effectively reverts to a more global transformation, resembling the behavior of appearance codes. This flexible, scale-dependent approach offers the best of both worlds, providing localized fine-tuning where necessary, while maintaining global consistency when appropriate.
We demonstrate that the multi-scale bilateral grid significantly enhances the geometric accuracy of dynamic, decoupled driving scenes, which is essential for autonomous driving applications where precise geometry plays a crucial role in tasks like obstacle avoidance and path planning. This is evidenced by our evaluation on multiple widely used datasets, including Waymo sun2020scalability , NuScenes caesar2020nuscenes , Argoverse wilson2023argoverse , and PandaSet xiao2021pandaset . The corresponding Chamfer Distance (CD) values, displayed in the Teaser. 1, highlight a marked reduction in error and improvement in reconstruction quality. Additionally, our method reduces photometric inconsistency, mitigating the appearance of floaters and enhancing the overall realism of the scene.
The contributions of this paper are as follows: First, we introduce a novel multi-scale bilateral grid that unifies appearance codes and bilateral grids, transitioning to either of the two paradigms in extreme cases, thus enhancing both modeling capability and optimization efficiency. Second, we show that by addressing photometric inconsistencies, it improves the geometric accuracy of dynamic, decoupled driving scene reconstructions. Third, we provide extensive benchmarking across four widely used datasets—Waymo, NuScenes, Argoverse, and PandaSet—where our method outperforms previous approaches, showcasing notable improvements in both qualitative and quantitative results.
Figure 2: Overview of our method. We unify appearance codes with multi-scale bilateral grids. Initially, a coarse rendering is obtained from a Gaussian scene graph. This rendered image is then processed by our multi-scale bilateral grids to perform detailed per-pixel color modeling, guided by a luminance-based map through slice and fusion operations.
2 Related Works
2.1 Bilateral Grids and Appearance Codes
Recent advancements in bilateral grids and appearance codes have significantly influenced techniques for addressing photometric inconsistencies, which are particularly critical in dynamic environments such as autonomous driving. The bilateral grids, first introduced for real-time edge-aware image processing chen2007real , enable efficient manipulation of spatial and intensity variations, forming the foundation for many modern techniques gharbi2017deep ; marki2016bilateral ; chen2016bilateral ; xu2021bilateral ; zheng2021ultra . Extensions like deep bilateral learning gharbi2017deep further enhance this framework by predicting local affine transformations in bilateral space, achieving real-time image enhancement on resource-constrained devices. Similarly, bilateral-guided upsampling chen2016bilateral and cost volume refinement xu2021bilateral have proven effective in tasks ranging from high-dynamic-range imaging to stereo matching, demonstrating the versatility of bilateral representations. In the context of neural rendering, methods such as NeRF tancik2022block ; jang2021codenerf ; martin2021nerf and Gaussian Splatting (GS) zhang2024gaussian ; kulhanek2024wildgaussians have achieved groundbreaking results in novel view synthesis but photometric inconsistency caused by varying illumination or transient occluders martin2021nerf ; zhang2024gaussian remains a challenge. Techniques like bilateral guided radiance field processing wang2024bilateral address these issues by disentangling camera-specific enhancements and reapplying them consistently in 3D space. WildGaussians kulhanek2024wildgaussians integrates DINO-based appearance codes into 3D GS to robustly handle occlusions and dynamic lighting in uncontrolled scenes. Cross-Ray NeRF yang2023cross employs cross-ray feature covariance and grid-based transient masking to harmonize appearance variations and suppress occlusions in unconstrained image collections. Despite these strides, existing approaches often struggle to balance global consistency with localized adaptability. For instance, while appearance codes provide global adjustments, they lack the granularity to model fine-grained variations durand2002fast . Conversely, bilateral grids excel at pixel-wise transformations but are challenging to optimize effectively paris2006fast . To address this issue, we introduce a multi-scale bilateral grid which unifies bilateral grids and appearance codes, facilitating high-quality dynamic autonomous driving scene reconstruction with enhanced modeling capability and optimization efficiency.
2.2 Autonomous Driving Simulation
Autonomous driving simulation has emerged as a critical tool for developing perception, planning, and control systems by generating diverse, realistic driving scenarios wu2023mars ; okae2021robust ; yan2024street ; lindstrom2024nerfs ; tao2024alignmif ; li2024uniscene . Recent advancements focus on photorealistic rendering, dynamic scene modeling, and multi-modal sensor simulation wu2023mars ; li2024hierarchical ; zhou2024drivinggaussian ; tonderski2024neurad ; yan2024street . For instance, MARS wu2023mars employs a modular NeRF-based framework to independently control static and dynamic scene elements, while Street Gaussian yan2024street achieves real-time urban scene rendering using explicit 3D Gaussian representations. DrivingGaussian zhou2024drivinggaussian enhances dynamic scene reconstruction via composite Gaussian Splatting, ensuring occlusion accuracy and multi-camera consistency. NeuRAD tonderski2024neurad integrates sensor-specific effects (e.g., rolling shutter, LiDAR beam divergence) to improve novel view synthesis. Holistic scene understanding and editing have also seen progress. HUGS zhou2024hugs combines static and dynamic 3D Gaussians for real-time semantic parsing. ChatSim wei2024editable enables language-driven scene editing with external asset integration. LiDAR data integration has advanced through methods like LiDAR-NeRF tao2024lidar , which uses structural regularization for low-texture regions. HO-Gaussian li2024ho , which merges grid-based volumes with Gaussian Splatting to eliminate Structure-from-Motion (SFM) dependencies.
Despite progress, gaps persist between simulated and real-world data lindstrom2024nerfs ; tao2024alignmif . AlignMiF tao2024alignmif addresses LiDAR-camera misalignments via geometry-aligned implicit fields, while RodUS nguyen2024rodus decomposes urban scenes into static and dynamic components using 4D semantics to reduce artifacts. Innovations in large-scale generation include InfiniCube lu2024infinicube , which leverages sparse-voxel representations for unbounded dynamic scenes, and Omnire chen2024omnire , which reconstructs diverse dynamic objects (e.g., pedestrians) for human-vehicle interaction simulations. Furthermore, recent works like GaussianPro cheng2024gaussianpro and SplatAD hess2024splatad explore to refine 3D GS for real-time rendering of dynamic scenes. GaussianPro cheng2024gaussianpro introduces progressive propagation for texture-less surfaces, and SplatAD hess2024splatad models sensor-specific phenomena (e.g., rolling shutter, LiDAR intensity). In this work, we aim to reconstruct high-quality driving scenes by unifying appearance codes and bilateral grids to enhance photometric consistency and modeling capability.
3 Methodology
3.1 Problem Formulation
We aim to reconstruct a 3D scene representation $\mathbb{G}$ from multi-view images and LiDAR depth maps, commonly available in autonomous driving datasets. Given a set of images $\{I_{c,t}\}$ and corresponding depths $\{D_{c,t}\}$ captured by $N$ cameras over $T$ time steps, we formulate our objective as minimizing the discrepancy between rendered and observed images and depths:
$\min_{\mathbb{G}}\sum_{c,t}(\left\|I_{c,t}^{r}-I_{c,t}\right\|_{2}^{2}+\lambda_{D}\left\|D_{c,t}^{r}-D_{c,t}\right\|_{2}^{2})~{}~{},$ (1)
where $I_{c,t}^{r}$ and $D_{c,t}^{r}$ are the rendered appearance and geometry of the scene $\mathbb{G}$ from camera $c$ at time $t$, and $\lambda_{D}$ is a weight balancing the depth term.
However, dynamic driving scenes and varying camera properties cause inconsistent appearances, leading to geometric and texture artifacts. To address inconsistency, we decompose each image $I_{c,t}$ into consistent $\mathcal{C}_{c,t}$ (e.g., intrinsic scene colors and constant sensor adjustments like normalization exposure), and non-consistent $\mathcal{N}_{c,t}$ (e.g., components-varying lighting conditions, different camera settings and diverse image ISP effects):
$I_{c,t}=\mathcal{C}_{c,t}+\mathcal{N}_{c,t}~{}~{},$ (2)
We model $\mathcal{N}_{c,t}$ as a non-linear transformation $\mathcal{F}(\mathcal{C}_{c,t})$. Substituting into Eq. (2):
$I_{c,t}=\mathcal{C}_{c,t}+\mathcal{F}(\mathcal{C}_{c,t})~{}~{},$ (3)
Let $\mathcal{E}(\mathcal{C}_{c,t})=\mathcal{C}_{c,t}+\mathcal{F}(\mathcal{C}_{c,t})$, we can reformulate the optimization objective in Eq. (1) as:
$\min_{\mathbb{G}}\sum_{c,t}(\left\|\mathcal{E}(\mathcal{C}_{c,t}^{r})-I_{c,t}\right\|_{2}^{2}+\lambda_{D}\left\|D_{c,t}^{r}-D_{c,t}\right\|_{2}^{2})~{}~{},$ (4)
This reformulation links photometric and geometric consistency optimization within a joint objective, guiding $\mathcal{C}_{c,t}^{r}$ adjustments by geometric constraints ($D_{c,t}^{r}$). This mitigates texture-geometry ambiguities (e.g., shadows on road surfaces misinterpretable as geometry changes). $\mathcal{F}(\mathcal{C}_{c,t})$ models transient appearance, while $\mathcal{C}_{c,t}^{r}$ enforces consistent scene properties. Experiments (Sec. 4, Tab. 1) show this joint optimization reduces both photometric error and geometric drift in comparison to baselines.
3.2 Gaussian Splatting for Autonomous Driving Environments
Following prior methodologies chen2024omnire ; tonderski2024neurad , we represent autonomous driving environments using a hybrid scene graph, decomposing the scene into sky, background, and dynamic object models. Each dynamic object is represented by a 3D model in canonical space and transformed to the scene with an associated sequence of SE(3) transformations. We derive the transformation matrixes from existing object detection pipelines or ground-truth annotations.
For the sky, we use an environment map to model sky color based on viewing direction, while static background is represented as a set of semi-transparent 3D Gaussians. Each Gaussian is characterized by a learnable opacity parameter $o\in(0,1)$, a mean position $\mu\in\mathbb{R}^{3}$, and an anisotropic covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$ which is parameterized by a scale vector $S\in\mathbb{R}^{3}$ and a rotation quaternion $q\in\mathbb{R}^{4}$. Additionally, Spherical harmonics coefficients $c\in\mathbb{R}^{F}$ are used to model appearance.
For the moving object models, we further distinguish between non-deformable (e.g., vehicles) and deformable objects (e.g., pedestrians). Non-deformable objects $\{\mathbb{G}_{i}^{N}|i\in\{1,\dots,n_{n}\}\}$ are optimized in their local coordinate space and transformed into the global world space via their pose $T$:
$\mathbb{G}_{i}^{N}=T\otimes\hat{\mathbb{G}}_{i}^{N}~{}~{},$ (5)
For deformable objects $\{\mathbb{G}_{i}^{D}|i\in\{1,\dots,n_{d}\}\}$, we employ a deformation network $\mathcal{F}_{\Phi}$ (parameterized by $\Phi$) to adapt the Gaussian representation based on latent variable $e$ and time $t$:
$\mathbb{G}_{i}^{D}=T\otimes\left(\hat{\mathbb{G}}_{i}^{D}\oplus\mathcal{F}_{\Phi}(\hat{\mathbb{G}}_{i}^{D},e,t)\right)~{}~{}.$ (6)
3.3 Multi-Scale Bilateral Grid for Appearance Enhancement
Solving the photometric correction problem formulated in Eq. (4) fundamentally depends on the ability to effectively model the complex and spatially-varying photometric transformations present in real-world driving scenes. The function $\mathcal{E}(\cdot)$ in our formulation represents this crucial photometric enhancement process. To address this, we propose the Multi-Scale Bilateral Grid architecture, detailed in this section, as a novel grounded solution for approximating $\mathcal{E}(\cdot)$ and achieving high-quality, consistent appearance enhancement.
3.3.1 Multi-Scale Bilateral Grid Architecture
To address limitations of existing methods in handling diverse photometric variations in driving scenes, we propose a multi-scale bilateral grid framework. Our framework achieves this unification by employing a hierarchical pyramid of bilateral grids, organized across multiple scales. This multi-scale design is directly inspired by the nature of photometric inconsistencies in real-world environments, which range from global scene-level changes to highly localized variations (e.g., from overall lighting shifts to fine texture-level shadows). By utilizing this grid hierarchy, our framework aims to capture and effectively correct photometric variations at their corresponding scales, thus enabling a more comprehensive and spatially adaptable photometric correction.
The multi-scale bilateral grid transformation can be formulated as $I^{e}=\bar{A}\odot I^{r}$, where $\bar{A}$ represents a composite, scale-dependent photometric transformation. Crucially, this composite transformation is constructed hierarchically, combining transformations learned by individual bilateral grids at different scales, progressing from coarse to fine. This staged composition is key to achieving scale-dependent adaptation to photometric inconsistencies and allows us to move beyond the inherent limitations of single-scale methods. Further details are provided in the subsequent sections.
Our framework achieves this hierarchical representation using a three-level bilateral grid pyramid (Fig. 2): (1) Coarse Level ($2\times 2\times 1\times 12$ grid) captures global scene structure and approximate global appearance codes; (2) Intermediate Level ($4\times 4\times 2\times 12$ grid) represents regional features and capture mid-range photometric variations; (3) Fine Level ($8\times 8\times 4\times 12$ grid) encodes local details and approximate pixel-wise bilateral grids within the multi-scale framework.
At each level $l$, the grid tensor is defined as $\mathcal{A}^{(l)}\in\mathbb{R}^{H^{(l)}\times W^{(l)}\times D^{(l)}\times C^{(l)}}$, where Spatial Dimension $(H^{(l)},W^{(l)})$ represents spatial resolution (height, width) at level $l$, Guidance Dimension $D^{(l)}$ specifies the guidance intensity levels at level $l$ and Coefficient Channel $C=12$ represents a flattened $3\times 4$ affine color transformation matrix.
3.3.2 Guidance Map, Slice Operation, and Multi-Scale Fusion
To achieve adaptive and spatially-varying photometric correction, our multi-scale bilateral grid framework employs a guidance map and a slicing operation chen2007real to retrieve localized transformations, followed by a hierarchical fusion strategy to combine transformations across scales.
We first derive a luminance-based guidance map $I^{g}(u,v)=\text{GrayScale}(I^{r}(u,v))$ from the rendered image $I^{r}(u,v)$, following the approach of chen2007real ; wang2024bilateral ; gharbi2017deep . This map encodes spatial brightness variations like shadows and highlights, serving as a spatially-varying query to our multi-scale grid. For each level $l$ and pixel $(u,v)$, the luminance $d$ from $I^{g}$ is used to perform a slicing operation, querying the grid’s intensity dimension $D^{(l)}$ to locate an intensity bin. Affine transformation coefficients $\mathcal{A}^{(l)}_{i,j,k}$ around this bin are then combined via trilinear interpolation for a level-specific transformation $\bar{A}^{(l)}(u,v)$. The slicing operation retrieves per-pixel transformations from each grid level, written as:
$\bar{A}^{(l)}(u,v)=\sum\limits_{i,j,k}w_{i,j,k}(u,v,d)\mathcal{A}^{s}(i,j,k)~{}~{},$ (7)
To efficiently fuse transformations across scales, we utilize hierarchical function composition. A naive approach of full-resolution slicing at each scale is computationally expensive and redundant due to local photometric coherence. To address this, we employ downsampled guidance maps $I^{g{(l)}}$ for each scale $l$, performing slicing to obtain low-resolution coefficient fields, which are then upsampled. This approach explicitly links scale-aware guidance to patch operations-the reduced-resolution maps enable efficient spatial aggregation of photometric transformations while enhances efficiency (See ablation study in Tab. 5). The level-specific transformations $\mathcal{T}^{(l)}$, decomposed into linear matrices $\bar{M}^{(l)}$ and translations $\bar{T}^{(l)}$, are then sequentially composed from coarse to fine to produce the final composite transformation $\bar{A}$. The refined image $I^{e}(u,v)$ is obtained by applying $\bar{A}(u,v)$ to $I^{r}(u,v)$, with the fusion process expressed as:
$I^{e}=\mathcal{T}^{(L-1)}\circ\mathcal{T}^{(L-2)}\circ\cdots\circ\mathcal{T}^{(0)}(I^{r})~{}~{},$ (8)
This hierarchical fusion decomposes the photometric transformation into scale-dependent residual refinements. Coarse scales capture global transformations, intermediate scales refine regional variations, and fine scales address local details (Further elaboration in Sec. 4.3). Further details on the interpolation kernel and computational efficiency considerations are provided in Appendix A1.1.
Dataset Method Full Image Human Vehicle Geometry
PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ CD $\downarrow$ RMSE $\downarrow$ Depth $\downarrow$
Waymo sun2020scalability Baseline 28.92 0.833 25.24 0.681 23.50 0.697 1.482 2.785 0.540
Baseline w/AC 28.95 0.835 24.92 0.656 23.51 0.699 1.378 2.760 0.519
Baseline w/BG 28.19 0.831 24.79 0.664 23.09 0.693 1.523 2.798 0.492
Ours 29.23 0.836 25.55 0.683 23.50 0.707 0.989 2.744 0.477
NuScenes caesar2020nuscenes Baseline 26.37 0.837 26.17 0.774 26.49 0.827 1.458 3.420 0.110
Baseline w/AC 26.38 0.840 25.83 0.762 26.52 0.828 1.437 3.415 0.106
Baseline w/BG 25.98 0.837 26.00 0.771 26.10 0.820 1.380 3.390 0.097
Ours 27.69 0.847 26.53 0.777 27.31 0.834 1.161 3.340 0.059
Argoverse wilson2023argoverse Baseline 24.59 0.848 23.35 0.743 24.47 0.812 0.954 4.208 0.050
Baseline w/AC 24.58 0.848 23.22 0.738 24.44 0.811 0.959 4.215 0.051
Baseline w/BG 23.31 0.842 22.99 0.730 23.49 0.795 0.901 4.218 0.049
Ours 24.68 0.849 23.35 0.739 24.51 0.827 0.807 4.199 0.040
Pandaset xiao2021pandaset Baseline 30.20 0.903 29.79 0.830 23.27 0.776 0.503 2.874 0.018
Baseline w/AC 30.20 0.903 29.64 0.824 23.39 0.783 0.496 2.871 0.020
Baseline w/BG 29.73 0.904 29.90 0.812 22.22 0.763 0.484 2.867 0.013
Ours 30.75 0.906 30.18 0.830 23.49 0.784 0.453 2.852 0.011
Table 1: Comparison with SOTA for scene reconstruction. w/AC represents appearance codes while w/BG represents single bilateral grids.
4 Experiments
4.1 Experimental Setup
Datasets.  We evaluate on four autonomous driving datasets: Waymo sun2020scalability , NuScenes caesar2020nuscenes , Argoverse wilson2023argoverse , and PandaSet xiao2021pandaset , selected for their diversity in sensor configurations (LiDAR, camera specifications), environmental conditions (lighting, seasons), and geographic locations. Identical model hyperparameters are used across all datasets to test generalization capability. Technical specifications including camera counts and ego-vehicle view cropping details are provided in Appendix A2.
Baseline.  Our focus is on modeling appearance variance across viewpoints, we build upon the open-source works DriveStudio chen2024omnire , ChatSim wei2024editable , StreetGS yan2024street and test three approaches: appearance codes, standalone bilateral grids, and our multi-scale bilateral grids. This allows for direct comparison of how each method handles photometric inconsistencies across viewpoints.
Method Reconstruction Novel View Synthesis Geometry
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ CD $\downarrow$ RMSE $\downarrow$ Depth $\downarrow$
ChatSim wei2024editable 25.10 0.805 0.252 23.27 0.725 0.270 1.557 3.509 0.106
ChatSim (Ours) 27.04 0.819 0.231 24.67 0.735 0.249 1.236 3.412 0.053
StreetGS yan2024street 25.74 0.822 0.240 23.64 0.736 0.259 1.604 3.544 0.107
StreetGS (Ours) 27.90 0.836 0.219 25.10 0.747 0.238 1.272 3.458 0.055
Table 2: Comparison with different baseline approaches for scene reconstruction. Performance is reported as the average over the nuScenes caesar2020nuscenes scenarios.
Implementation details.  We evaluate our method on many autonomous driving datasets to demonstrate its robustness. For each dataset, We train all the methods on a single NVIDIA L40 GPU and compare the 3D reconstruction and novel view synthesis results between camera and LiDAR data. Additionally, we ablate key components of our method and quantify their impact on both appearance and geometry quality.
Dataset Method PSNR $\uparrow$
Full Image Human Vehicle
Waymo sun2020scalability Baseline 26.40 23.91 21.25
Baseline w/AC 26.44 23.81 21.34
Baseline w/BG 21.51 22.91 19.82
Ours 26.55 23.97 21.36
NuScenes caesar2020nuscenes Baseline 23.74 23.55 22.80
Baseline w/AC 23.74 23.35 22.84
Baseline w/BG 23.60 23.63 22.79
Ours 24.64 23.75 23.22
Argoverse wilson2023argoverse Baseline 22.53 20.73 20.84
Baseline w/AC 22.51 20.70 20.80
Baseline w/BG 21.70 20.67 20.59
Ours 22.58 20.79 20.86
Pandaset xiao2021pandaset Baseline 27.49 26.94 21.02
Baseline w/AC 27.51 26.98 21.13
Baseline w/BG 27.38 27.08 20.77
Ours 27.89 27.22 21.32
Table 3: Comparison on Novel View Synthesis.
Training and Dynamic Rendering.
1) Training.  To optimize our multi-scale Gaussian scene representation, we employ a joint training strategy that minimizes a composite reconstruction loss :
$\mathcal{L}_{recon}=\lambda_{r}\mathcal{L}_{1}+\lambda_{s}\mathcal{L}_{\text{SSIM}}+\lambda_{d}\mathcal{L}_{d}+\lambda_{o}\mathcal{L}_{o}~{}~{},$ (9)
Furthermore, we introduce two additional regularization terms to enhance image fidelity: Adaptive Total Variation Regularization. This term encourages smoothness and reduces noise while preserving image details, and Circle Regularization Loss. This loss applies an inverse transformation to the ground-truth images, preventing discrepancies and image quality degradation. As visually demonstrated in Fig. 5, these two terms significantly improve image fidelity. Detailed explanations are provided in Appendix A1.
2) Rendering. For dynamic rendering using our multi-scale bilateral grid framework, we employ a specific interpolation strategy when encountering novel test images, especially those simulating dynamic ISP conditions. Our proposed method first conducts a temporal proximity search to identify the most relevant grids. Following this, we perform scale-specific interpolation using the two nearest grids found. More comprehensive details regarding this interpolation approach are available in Appendix A1.
Figure 3: Visualizations of our method and baseline methods on Waymo sun2020scalability , NuScenes caesar2020nuscenes , Argoverse wilson2023argoverse , and PandaSet xiao2021pandaset dataset.
Figure 4: Our proposed framework (c) outperforms appearance codes (a) and single bilateral grids (b) by simultaneously addressing optimization challenges and enhancing geometric modeling. This delivers significant geometric accuracy improvements (lower Chamfer Distance (CD) via multi-scale bilateral grid, reducing error and improving reconstruction) and reduced photometric inconsistency (fewer floaters) in dynamic, decoupled driving scenes. Yellow indicates a high lidar error, while Purple indicates a low lidar error.
4.2 Quantitative Evaluation
Reconstruction Quality.  We evaluate the reconstruction quality of our method using appearance and geometry metrics, with detailed results presented in Tab. 1.
(1) Appearance Fidelity. Our approach demonstrates superior performance in appearance-based metrics (PSNR and SSIM). As shown in Tab. 1, our method achieves the highest PSNR and SSIM scores for full image reconstruction across all tested datasets (Waymo, NuScenes, Argoverse, and Pandaset). Furthermore, when evaluating segmented dynamic regions such as humans and vehicles, our method largely surpasses the baselines, indicating its robustness in capturing both global scene appearance and fine-grained details of dynamic objects. For instance, on NuScenes, our PSNR for ’Vehicle’ is 27.31, compared to the best baseline score of 26.52. While there are isolated cases where a baseline performs comparably or slightly better (e.g., Human SSIM on Argoverse), our method generally offers the most consistent and highest quality appearance reconstruction.
(2) Geometric Accuracy. In terms of geometric fidelity, our method consistently outperforms all baselines across all datasets, as evidenced by the Chamfer Distance (CD), Root Mean Square Error (RMSE), and Depth L2 Error metrics in Tab. 1. For example, on the Waymo dataset, our method achieves a CD of 0.989, significantly lower than the best baseline score of 1.378. This superior geometric accuracy is partly attributed to our model’s improved handling of "floater" artifacts, leading to more precise LiDAR representations, as qualitatively illustrated in Fig. 4.
Novel View Synthesis.  The performance of our method in novel view synthesis is detailed in Tab. 3. Our approach consistently achieves the highest PSNR scores across all datasets for full images as well as for segmented human and vehicle categories. For example, on the NuScenes dataset, our method achieves a PSNR of 24.64 for the full image, 23.75 for humans, and 23.22 for vehicles, outperforming all baseline variations. This underscores the capability of our method to generate high-fidelity novel views from new camera poses.
Comparison with State-of-the-Art Methods. To further assess the benefits of our architectural contributions, we integrated our core components into two recent state-of-the-art methods: ChatSim wei2024editable and StreetGS yan2024street . The comparative results, averaged over NuScenes scenarios, are presented in Tab. 2.
When our enhancements are applied (denoted as "ChatSim (Ours)" and "StreetGS (Ours)"), both methods exhibit significant improvements across all evaluated aspects. For instance, integrating our approach with ChatSim improves its reconstruction PSNR from 25.10 to 27.04 and its novel view synthesis PSNR from 23.27 to 24.67. Similarly, StreetGS sees its reconstruction PSNR boosted from 25.74 to 27.90 and its Chamfer Distance reduced from 1.604 to 1.272. These results highlight the generalizability and effectiveness of our proposed techniques, demonstrating their capability to enhance the performance of existing strong baseline models for comprehensive scene understanding and representation.
4.3 Qualitative Evaluation
Rendering Results.  Fig. 3 presents visualization results of our method on scenes from Waymo sun2020scalability , NuScenes caesar2020nuscenes , Argoverse wilson2023argoverse , and PandaSet xiao2021pandaset , demonstrating its capability to handle diverse conditions that might otherwise lead to inconsistencies.
Figure 5: Visualizations of ablations on Circle Regularization loss and Adaptive TV loss.
Figure 6: Visualization of Affine Transformations and Distributions. The first three rows display the affine transformation results and corresponding distributions when applying the first level (scale 0), the first two levels (scales 0-1), and all three levels (scales 0-1-2) of our multi-scale approach. The final row compares these to results from a single bilateral grid, demonstrating that the multi-scale bilateral grid achieves a coarse-to-fine, smooth image enhancement with more diverse processing capabilities.
Affine Transformations Distributions.  According to Fig. 6, analyzing grayscale histograms of learned affine transformations reveals a key distinction between original single-scale and our multi-scale bilateral grids. For the original grid, each viewpoint’s histogram typically exhibits two distinct peaks. This bi-modal distribution suggests a focused, somewhat limited photometric adaptation. It implies that for each view, correction focuses on a few dominant transformations. More visualization results are shown at Fig. 3. A particularly interesting observation arises examining histograms of our multi-scale grid’s coarsest level. Intriguingly, these histograms also demonstrate peaky distributions, varying significantly across viewpoints. This isn’t contradictory, but insightful: even at its coarsest level, the grid learns view-dependent base appearances. The coarse grid captures dominant, scene-wide photometric transformations characteristic of each viewpoint. Peaks in coarse-level histograms likely represent these primary, global photometric adjustments, tailored per view.
Configuration Grid Params (M) Training Time (h) Testing time (FPS)
Baseline w/AC 0.072 1.93 53
Baseline w/BG 27.843 2.85 54 / 38†
Ours 3.969 2.10 54 / 42†
Table 4: Computational Efficiency Analysis. †  means using bilateral grids processing during rendering.
However, the key advantage is evident in the overlaid histogram, which aggregates distributions from all multi-scale levels and viewpoints. This histogram is notably flatter and more dispersed than those of the original grid, a characteristic directly resulting from hierarchical and residual refinement. Our framework first uses the coarse grid to establish a view-dependent base appearance. Subsequently, medium and fine grids learn residual transformations to correct coarse grid errors. Medium grids then refine regional photometric variations, while fine grids handle local, pixel-level details. This layered, residual approach significantly broadens the spectrum of learnable photometric transformations. The original grid’s two-peaked histograms signify its narrow adaptation. In contrast, the multi-scale version’s flatter histogram reveals its core strength: a diverse, generalizable, and robust representation of spatially-varying photometric transformations. This enhanced representation is crucial for consistent and high-quality renderings across diverse driving views.
4.4 Ablation Studies
As shown in Tab. 4, Tab. 5 and Fig. 5, we systematically evaluate component contributions through:
Settings Recon.$\uparrow$ NVS.$\uparrow$ CD $\downarrow$
   Full model 27.69 24.64 1.161
(a) Guidance Map Resolution
   (1,1,1) 27.62 24.34 1.271
   (2,2,1) 27.68 24.64 1.170
   (8,8,4) 27.52 24.53 1.187
   (16,16,8) 27.50 24.54 1.202
(b) Grid Size Combinations
   (4,4,2)+(8,8,4) 27.49 24.56 1.963
   (2,2,1)+(8,8,4) 27.57 24.57 1.940
   (2,2,1)+(4,4,2) 27.33 24.42 1.213
   (4,4,2)+(8,8,4)+(16,16,8) 27.31 24.46 2.297
(c) Loss Functions
   w/o circle regularization loss 27.47 24.52 1.164
   w/o adaptive total variance loss 27.64 24.61 1.169
Table 5: Ablations. Numbers in Guidance Map Resolution settings indicate downsample factors from coarse level to fine level, and numbers in Grid Size Combinations settings indicate grid size.
Guidance Map Resolution.   Varying guidance map resolution shows a trade-off in performance. Using the highest resolution (1,1,1) results in a slightly lower NVS PSNR compared to using (2,2,1) which achieves the highest NVS PSNR(24.64) in these ablations, demonstrating our patch-based approach balances efficiency and effectiveness. Grid Size Combinations.   Using different combinations of grid sizes significantly impacts geometry, as indicated by CD. Combining smaller grids like (2,2,1)+(4,4,2) results in a lower CD (1.213) than using combinations including larger grids or the combination with the largest grid (4,4,2)+(8,8,4)+(16,16,8). Loss Functions.   Ablating the circle regularization loss slightly reduces both Reconstruction PSNR (from 27.69 to 27.47) and NVS PSNR (from 24.64 to 24.52). Removing the adaptive total variance loss has a minimal impact on rendering result. Efficiency.   As shown in Tab. 4, while the ’Baseline w/BG’ dramatically increases parameters and training time, our method achieves a much more moderate increase in parameters and training time (to 2.10 hours) relative to the lightweight appearance codes. With multi-scale bilateral grid guided training, we found the rendering speed may improve due to the decreased number of unnecessary Gaussians.
5 Conclusion
We propose a multi-scale bilateral grid seamlessly integrating global appearance codes and pixel-wise bilateral grids to address photometric inconsistencies in autonomous driving neural rendering. Limitations. While our proposed multi-scale bilateral grid demonstrates significant improvements in handling photometric inconsistencies and enhancing geometric accuracy in driving scene reconstruction, certain limitations warrant discussion. (1) The computational remains higher than simpler appearance code techniques; this could be a factor in highly resource-constrained deployment scenarios. (2) while effective for many dynamic elements, modeling extremely complex, fast-moving, or highly non-rigid objects with complete fidelity remains an ongoing challenge.

### v4

@@
 Neural rendering techniques, including NeRF and Gaussian Splatting (GS), rely on photometric consistency to produce high-quality reconstructions.
-However, in real-world scenarios, it is challenging to guarantee perfect photometric consistency in acquired images.
+However, in real-world driving scenarios, it is challenging to guarantee perfect photometric consistency in acquired images.
 Appearance codes have been widely used to address this issue, but their modeling capability is limited, as a single code is applied to the entire image.
@@
 Our code is open-sourced at https://bigcileng.github.io/bilateral-driving.
-Figure 1: (a) Appearance codes rely on global affine transformations but have limited modeling capability.
+Figure 1: Unifying appearance codes and bilateral grids.
+(a) Appearance codes rely on global affine transformations but have limited modeling capability.
 (b) Bilateral grids perform pixel-wise transformations, improving photometric consistency but are challenging to optimize.
-(c) The proposed multi-scale bilateral grid unifies appearance codes and bilateral grids, enabling patch-wise transformations.
+(c) The proposed multi-scale bilateral grid unifies both paradigms, enabling patch-wise transformations.
 1 Introduction
-Neural rendering techniques, such as NeRFs  mildenhall2021nerf ; wang2021neus ; yu2022monosdf ; liu2024rip ; yuan2024slimmerf and Gaussian Splatting (GS) kerbl20233d ; chen2024pgsr ; cheng2024gaussianpro ; zhang2024drone ; song2024sa , have demonstrated significant potential in producing high-quality 3D reconstructions by leveraging photometric consistency.
-However, ensuring photometric consistency across images in real-world scenarios remains a challenge, as lighting conditions, viewpoints, and camera settings can vary considerably martin2021nerf ; zhang2024gaussian ; kulhanek2024wildgaussians ; wang2025semantic .
-In autonomous driving applications yan2024renderworld ; wu2023mars ; tonderski2024neurad ; chen2024omnire ; li2024uniscene ; jin2024tod3cap ; tian2023unsupervised , these challenges are amplified due to the presence of multiple cameras capturing the same scene from different angles over time.
+Neural rendering techniques, such as NeRFs  Mildenhall et al.
+(2021); Wang et al.
+(2021); Yu et al.
+(2022); Liu et al.
+(2024); Yuan and Zhao (2024) and Gaussian Splatting (GS) Kerbl et al.
+(2023); Chen et al.
+(2024a); Cheng et al.
+(2024); Zhang et al.
+(2024b); Song et al.
+(2024); Ye et al.
+(2025); Shuai et al.
+(2025), have demonstrated significant potential in producing high-quality 3D reconstructions by leveraging photometric consistency.
+However, ensuring photometric consistency across images in real-world scenarios remains a challenge, as lighting conditions, viewpoints, and camera settings can vary considerably Martin-Brualla et al.
+(2021); Zhang et al.
+(2024a); Kulhanek et al.
+(2024); Wang et al.
+(2025); Gao et al.
+(2024).
+In autonomous driving applications Yan et al.
+(2024b); Wu et al.
+(2023); Tonderski et al.
+(2024); Chen et al.
+(2024b); Li et al.
+(2024b); Jin et al.
+(2024); Tian et al.
+(2023); Han et al.
+(2025); Yu et al.
+(2025), these challenges are amplified due to the presence of multiple cameras capturing the same scene from different angles over time.
 As illustrated in the left half of Teaser.
@@
 These inconsistencies pose difficulties in accurately reconstructing dynamic driving environments.
-To address this issue, appearance codes tonderski2024neurad ; martin2021nerf ; chen2024omnire have been introduced to encode per-image information and assist in correcting the photometric discrepancies.
+To address this issue, appearance codes Tonderski et al.
+(2024); Martin-Brualla et al.
+(2021); Chen et al.
+(2024b) have been introduced to encode per-image information and assist in correcting the photometric discrepancies.
 These codes, while effective, have limitations in modeling capability since they apply a single transformation across the entire image, often neglecting local variations that may occur in complex scenes.
@@
 1 NuScenes (a), the fence is blurred).
-Recent advancements wang2024bilateral ; gharbi2017deep have introduced bilateral grids to perform pixel-wise transformations, enabling improved photometric consistency by allowing more localized adjustments.
+Recent advancements Wang et al.
+(2024); Gharbi et al.
+(2017) have introduced bilateral grids to perform pixel-wise transformations, enabling improved photometric consistency by allowing more localized adjustments.
 However, these grids face significant challenges in optimization, as they require complex constraints to avoid instability during training (As shown in Teaser.
@@
 We demonstrate that the multi-scale bilateral grid significantly enhances the geometric accuracy of dynamic, decoupled driving scenes, which is essential for autonomous driving applications where precise geometry plays a crucial role in tasks like obstacle avoidance and path planning.
-This is evidenced by our evaluation on multiple widely used datasets, including Waymo sun2020scalability , NuScenes caesar2020nuscenes , Argoverse wilson2023argoverse , and PandaSet xiao2021pandaset .
+This is evidenced by our evaluation on multiple widely used datasets, including Waymo Sun et al.
+(2020), NuScenes Caesar et al.
+(2020), Argoverse Wilson et al.
+(2023), and PandaSet Xiao et al.
+(2021).
 The corresponding Chamfer Distance (CD) values, displayed in the Teaser.
@@
 We unify appearance codes with multi-scale bilateral grids.
-Initially, a coarse rendering is obtained from a Gaussian scene graph.
-This rendered image is then processed by our multi-scale bilateral grids to perform detailed per-pixel color modeling, guided by a luminance-based map through slice and fusion operations.
+A coarse rendering from the Gaussian scene graph is refined by multi-scale bilateral grids to model per-pixel color with a luminance-guided slice-and-fuse pipeline.
 2 Related Works
@@
 Recent advancements in bilateral grids and appearance codes have significantly influenced techniques for addressing photometric inconsistencies, which are particularly critical in dynamic environments such as autonomous driving.
-The bilateral grids, first introduced for real-time edge-aware image processing chen2007real , enable efficient manipulation of spatial and intensity variations, forming the foundation for many modern techniques gharbi2017deep ; marki2016bilateral ; chen2016bilateral ; xu2021bilateral ; zheng2021ultra .
-Extensions like deep bilateral learning gharbi2017deep further enhance this framework by predicting local affine transformations in bilateral space, achieving real-time image enhancement on resource-constrained devices.
-Similarly, bilateral-guided upsampling chen2016bilateral and cost volume refinement xu2021bilateral have proven effective in tasks ranging from high-dynamic-range imaging to stereo matching, demonstrating the versatility of bilateral representations.
-In the context of neural rendering, methods such as NeRF tancik2022block ; jang2021codenerf ; martin2021nerf and Gaussian Splatting (GS) zhang2024gaussian ; kulhanek2024wildgaussians have achieved groundbreaking results in novel view synthesis but photometric inconsistency caused by varying illumination or transient occluders martin2021nerf ; zhang2024gaussian remains a challenge.
-Techniques like bilateral guided radiance field processing wang2024bilateral address these issues by disentangling camera-specific enhancements and reapplying them consistently in 3D space.
-WildGaussians kulhanek2024wildgaussians integrates DINO-based appearance codes into 3D GS to robustly handle occlusions and dynamic lighting in uncontrolled scenes.
-Cross-Ray NeRF yang2023cross employs cross-ray feature covariance and grid-based transient masking to harmonize appearance variations and suppress occlusions in unconstrained image collections.
+The bilateral grids, first introduced for real-time edge-aware image processing Chen et al.
+(2007), enable efficient manipulation of spatial and intensity variations, forming the foundation for many modern techniques Gharbi et al.
+(2017); Märki et al.
+(2016); Chen et al.
+(2016); Xu et al.
+(2021); Zheng et al.
+(2021).
+Extensions like deep bilateral learning Gharbi et al.
+(2017) further enhance this framework by predicting local affine transformations in bilateral space, achieving real-time image enhancement on resource-constrained devices.
+Similarly, bilateral-guided upsampling Chen et al.
+(2016) and cost volume refinement Xu et al.
+(2021) have proven effective in tasks ranging from high-dynamic-range imaging to stereo matching, demonstrating the versatility of bilateral representations.
+In the context of neural rendering, methods such as NeRF Tancik et al.
+(2022); Jang and Agapito (2021); Martin-Brualla et al.
+(2021) and Gaussian Splatting (GS) Zhang et al.
+(2024a); Kulhanek et al.
+(2024) have achieved groundbreaking results in novel view synthesis but photometric inconsistency caused by varying illumination or transient occluders Martin-Brualla et al.
+(2021); Zhang et al.
+(2024a) remains a challenge.
+Techniques like bilateral guided radiance field processing Wang et al.
+(2024) address these issues by disentangling camera-specific enhancements and reapplying them consistently in 3D space.
+WildGaussians Kulhanek et al.
+(2024) integrates DINO-based appearance codes into 3D GS to robustly handle occlusions and dynamic lighting in uncontrolled scenes.
+Cross-Ray NeRF Yang et al.
+(2023b) employs cross-ray feature covariance and grid-based transient masking to harmonize appearance variations and suppress occlusions in unconstrained image collections.
 Despite these strides, existing approaches often struggle to balance global consistency with localized adaptability.
-For instance, while appearance codes provide global adjustments, they lack the granularity to model fine-grained variations durand2002fast .
-Conversely, bilateral grids excel at pixel-wise transformations but are challenging to optimize effectively paris2006fast .
+For instance, while appearance codes provide global adjustments, they lack the granularity to model fine-grained variations Durand and Dorsey (2002).
+Conversely, bilateral grids excel at pixel-wise transformations but are challenging to optimize effectively Paris and Durand (2006).
 To address this issue, we introduce a multi-scale bilateral grid which unifies bilateral grids and appearance codes, facilitating high-quality dynamic autonomous driving scene reconstruction with enhanced modeling capability and optimization efficiency.
 2.2 Autonomous Driving Simulation
-Autonomous driving simulation has emerged as a critical tool for developing perception, planning, and control systems by generating diverse, realistic driving scenarios wu2023mars ; okae2021robust ; yan2024street ; lindstrom2024nerfs ; tao2024alignmif ; li2024uniscene .
-Recent advancements focus on photorealistic rendering, dynamic scene modeling, and multi-modal sensor simulation wu2023mars ; li2024hierarchical ; zhou2024drivinggaussian ; tonderski2024neurad ; yan2024street .
-For instance, MARS wu2023mars employs a modular NeRF-based framework to independently control static and dynamic scene elements, while Street Gaussian yan2024street achieves real-time urban scene rendering using explicit 3D Gaussian representations.
-DrivingGaussian zhou2024drivinggaussian enhances dynamic scene reconstruction via composite Gaussian Splatting, ensuring occlusion accuracy and multi-camera consistency.
-NeuRAD tonderski2024neurad integrates sensor-specific effects (e.g., rolling shutter, LiDAR beam divergence) to improve novel view synthesis.
+Autonomous driving simulation has emerged as a critical tool for developing perception, planning, and control systems by generating diverse, realistic driving scenarios Wu et al.
+(2023); Okae et al.
+(2021); Yan et al.
+(2024a); Lindström et al.
+(2024); Tao et al.
+(2024b); Li et al.
+(2024b); Yang et al.
+(2024a); Gao et al.
+(2023).
+Recent advancements focus on photorealistic rendering, dynamic scene modeling, and multi-modal sensor simulation Wu et al.
+(2023); Li et al.
+(2024a); Zhou et al.
+(2024b); Tonderski et al.
+(2024); Yan et al.
+(2024a); Yang et al.
+(2023a, 2025b); Xiao et al.
+(2025).
+For instance, MARS Wu et al.
+(2023) employs a modular NeRF-based framework to independently control static and dynamic scene elements, while Street Gaussian Yan et al.
+(2024a) achieves real-time urban scene rendering using explicit 3D Gaussian representations.
+DrivingGaussian Zhou et al.
+(2024b) enhances dynamic scene reconstruction via composite Gaussian Splatting, ensuring occlusion accuracy and multi-camera consistency.
+NeuRAD Tonderski et al.
+(2024) integrates sensor-specific effects (e.g., rolling shutter, LiDAR beam divergence) to improve novel view synthesis.
 Holistic scene understanding and editing have also seen progress.
-HUGS zhou2024hugs combines static and dynamic 3D Gaussians for real-time semantic parsing.
-ChatSim wei2024editable enables language-driven scene editing with external asset integration.
-LiDAR data integration has advanced through methods like LiDAR-NeRF tao2024lidar , which uses structural regularization for low-texture regions.
-HO-Gaussian li2024ho , which merges grid-based volumes with Gaussian Splatting to eliminate Structure-from-Motion (SFM) dependencies.
-Despite progress, gaps persist between simulated and real-world data lindstrom2024nerfs ; tao2024alignmif .
-AlignMiF tao2024alignmif addresses LiDAR-camera misalignments via geometry-aligned implicit fields, while RodUS nguyen2024rodus decomposes urban scenes into static and dynamic components using 4D semantics to reduce artifacts.
-Innovations in large-scale generation include InfiniCube lu2024infinicube , which leverages sparse-voxel representations for unbounded dynamic scenes, and Omnire chen2024omnire , which reconstructs diverse dynamic objects (e.g., pedestrians) for human-vehicle interaction simulations.
-Furthermore, recent works like GaussianPro cheng2024gaussianpro and SplatAD hess2024splatad explore to refine 3D GS for real-time rendering of dynamic scenes.
-GaussianPro cheng2024gaussianpro introduces progressive propagation for texture-less surfaces, and SplatAD hess2024splatad models sensor-specific phenomena (e.g., rolling shutter, LiDAR intensity).
+HUGS Zhou et al.
+(2024a) combines static and dynamic 3D Gaussians for real-time semantic parsing.
+ChatSim Wei et al.
+(2024) enables language-driven scene editing with external asset integration.
+LiDAR data integration has advanced through methods like LiDAR-NeRF Tao et al.
+(2024a), which uses structural regularization for low-texture regions.
+HO-Gaussian Li et al.
+(2024c) merges grid-based volumes with Gaussian Splatting to eliminate Structure-from-Motion dependencies.
+Despite progress, gaps persist between simulated and real-world data Lindström et al.
+(2024); Tao et al.
+(2024b); Yang et al.
+(2024b, 2025a).
+AlignMiF Tao et al.
+(2024b) addresses LiDAR-camera misalignments via geometry-aligned implicit fields, while RodUS Nguyen et al.
+(2024) decomposes urban scenes into static and dynamic components using 4D semantics to reduce artifacts.
+Innovations in large-scale generation include InfiniCube Lu et al.
+(2024), which leverages sparse-voxel representations for unbounded dynamic scenes, and Omnire Chen et al.
+(2024b), which reconstructs diverse dynamic objects (e.g., pedestrians) for human-vehicle interaction simulations.
+Furthermore, recent works like GaussianPro Cheng et al.
+(2024) and SplatAD Hess et al.
+(2024) explore to refine 3D GS for real-time rendering of dynamic scenes.
+GaussianPro Cheng et al.
+(2024) introduces progressive propagation for texture-less surfaces, and SplatAD Hess et al.
+(2024) models sensor-specific phenomena (e.g., rolling shutter, LiDAR intensity).
 In this work, we aim to reconstruct high-quality driving scenes by unifying appearance codes and bilateral grids to enhance photometric consistency and modeling capability.
@@
 3.2 Gaussian Splatting for Autonomous Driving Environments
-Following prior methodologies chen2024omnire ; tonderski2024neurad , we represent autonomous driving environments using a hybrid scene graph, decomposing the scene into sky, background, and dynamic object models.
+Following prior methodologies Chen et al.
+(2024b); Tonderski et al.
+(2024), we represent autonomous driving environments using a hybrid scene graph, decomposing the scene into sky, background, and dynamic object models.
 Each dynamic object is represented by a 3D model in canonical space and transformed to the scene with an associated sequence of SE(3) transformations.
@@
 3.3.2 Guidance Map, Slice Operation, and Multi-Scale Fusion
-To achieve adaptive and spatially-varying photometric correction, our multi-scale bilateral grid framework employs a guidance map and a slicing operation chen2007real to retrieve localized transformations, followed by a hierarchical fusion strategy to combine transformations across scales.
-We first derive a luminance-based guidance map $I^{g}(u,v)=\text{GrayScale}(I^{r}(u,v))$ from the rendered image $I^{r}(u,v)$, following the approach of chen2007real ; wang2024bilateral ; gharbi2017deep .
+To achieve adaptive and spatially-varying photometric correction, our multi-scale bilateral grid framework employs a guidance map and a slicing operation Chen et al.
+(2007) to retrieve localized transformations, followed by a hierarchical fusion strategy to combine transformations across scales.
+We first derive a luminance-based guidance map $I^{g}(u,v)=\text{GrayScale}(I^{r}(u,v))$ from the rendered image $I^{r}(u,v)$, following the approach of Chen et al.
+(2007); Wang et al.
+(2024); Gharbi et al.
+(2017).
 This map encodes spatial brightness variations like shadows and highlights, serving as a spatially-varying query to our multi-scale grid.
@@
 The slicing operation retrieves per-pixel transformations from each grid level, written as:
-$\bar{A}^{(l)}(u,v)=\sum\limits_{i,j,k}w_{i,j,k}(u,v,d)\mathcal{A}^{s}(i,j,k)~{}~{},$ (7)
+$\bar{A}^{(l)}(u,v)=\sum\limits_{i,j,k}w_{i,j,k}(u,v,d)\mathcal{A}^{(l)}(i,j,k)~~,$ (7)
 To efficiently fuse transformations across scales, we utilize hierarchical function composition.
@@
 Further details on the interpolation kernel and computational efficiency considerations are provided in Appendix A1.1.
-Dataset Method Full Image Human Vehicle Geometry
-PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ CD $\downarrow$ RMSE $\downarrow$ Depth $\downarrow$
-Waymo sun2020scalability Baseline 28.92 0.833 25.24 0.681 23.50 0.697 1.482 2.785 0.540
-Baseline w/AC 28.95 0.835 24.92 0.656 23.51 0.699 1.378 2.760 0.519
-Baseline w/BG 28.19 0.831 24.79 0.664 23.09 0.693 1.523 2.798 0.492
-Ours 29.23 0.836 25.55 0.683 23.50 0.707 0.989 2.744 0.477
-NuScenes caesar2020nuscenes Baseline 26.37 0.837 26.17 0.774 26.49 0.827 1.458 3.420 0.110
-Baseline w/AC 26.38 0.840 25.83 0.762 26.52 0.828 1.437 3.415 0.106
-Baseline w/BG 25.98 0.837 26.00 0.771 26.10 0.820 1.380 3.390 0.097
-Ours 27.69 0.847 26.53 0.777 27.31 0.834 1.161 3.340 0.059
-Argoverse wilson2023argoverse Baseline 24.59 0.848 23.35 0.743 24.47 0.812 0.954 4.208 0.050
-Baseline w/AC 24.58 0.848 23.22 0.738 24.44 0.811 0.959 4.215 0.051
-Baseline w/BG 23.31 0.842 22.99 0.730 23.49 0.795 0.901 4.218 0.049
-Ours 24.68 0.849 23.35 0.739 24.51 0.827 0.807 4.199 0.040
-Pandaset xiao2021pandaset Baseline 30.20 0.903 29.79 0.830 23.27 0.776 0.503 2.874 0.018
-Baseline w/AC 30.20 0.903 29.64 0.824 23.39 0.783 0.496 2.871 0.020
-Baseline w/BG 29.73 0.904 29.90 0.812 22.22 0.763 0.484 2.867 0.013
-Ours 30.75 0.906 30.18 0.830 23.49 0.784 0.453 2.852 0.011
-Table 1: Comparison with SOTA for scene reconstruction. w/AC represents appearance codes while w/BG represents single bilateral grids.
+Table 1: Ablation studies and comparisons on four large-scale driving datasets.
+Our proposed method consistently outperforms OmniRe across all metrics and datasets.
+w/AC denotes appearance codes; w/BG denotes single bilateral grids (size of $16\times 16\times 8$).
+Dataset Method Reconstruction Novel View Synthesis Geometry
+PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ CD $\downarrow$ RMSE $\downarrow$ Depth $\downarrow$
+Waymo OmniRe $28.92$ $0.833$ 0.295 $26.40$ 0.761 0.311 $1.482$ $2.785$ $0.540$
+OmniRe w/AC $28.95$ $0.835$ 0.293 $26.44$ 0.759 0.315 $1.378$ $2.760$ $0.519$
+OmniRe w/BG $28.19$ $0.831$ 0.292 $21.51$ 0.743 0.333 $1.523$ $2.798$ $0.492$
+Ours 29.23 0.836 0.289 26.55 0.762 0.310 0.989 2.744 0.477
+NuScenes OmniRe $26.37$ $0.837$ 0.209 $23.74$ 0.733 0.232 $1.458$ $3.420$ $0.110$
+OmniRe w/AC $26.38$ $0.840$ 0.204 $23.74$ 0.732 0.229 $1.437$ $3.415$ $0.106$
+OmniRe w/BG $25.98$ $0.837$ 0.209 $23.60$ 0.705 0.262 $1.380$ $3.390$ $0.097$
+Ours 27.69 0.847 0.193 24.64 0.739 0.216 1.161 3.340 0.059
+Argoverse OmniRe $24.59$ $0.848$ 0.202 $22.53$ 0.755 0.220 $0.954$ $4.208$ $0.050$
+OmniRe w/AC $24.58$ $0.848$ 0.201 $22.51$ 0.756 0.219 $0.959$ $4.215$ $0.051$
+OmniRe w/BG $23.31$ $0.842$ 0.216 $21.70$ 0.725 0.254 $0.901$ $4.218$ $0.049$
+Ours 24.68 0.849 0.200 22.58 0.756 0.217 0.807 4.199 0.040
+PandaSet OmniRe $30.20$ $0.903$ 0.219 $27.49$ 0.835 0.240 $0.503$ $2.874$ $0.018$
+OmniRe w/AC $30.20$ $0.903$ 0.220 $27.51$ 0.841 0.242 $0.496$ $2.871$ $0.020$
+OmniRe w/BG $29.73$ $0.904$ 0.220 $27.38$ 0.830 0.246 $0.484$ $2.867$ $0.013$
+Ours 30.75 0.906 0.213 27.89 0.847 0.235 0.453 2.852 0.011
 4 Experiments
@@
 Datasets.
-We evaluate on four autonomous driving datasets: Waymo sun2020scalability , NuScenes caesar2020nuscenes , Argoverse wilson2023argoverse , and PandaSet xiao2021pandaset , selected for their diversity in sensor configurations (LiDAR, camera specifications), environmental conditions (lighting, seasons), and geographic locations.
+We evaluate on four autonomous driving datasets: Waymo Sun et al.
+(2020), NuScenes Caesar et al.
+(2020), Argoverse Wilson et al.
+(2023), and PandaSet Xiao et al.
+(2021), selected for their diversity in sensor configurations (LiDAR, camera specifications), environmental conditions (lighting, seasons), and geographic locations.
 Identical model hyperparameters are used across all datasets to test generalization capability.
@@
 Baseline.
-Our focus is on modeling appearance variance across viewpoints, we build upon the open-source works DriveStudio chen2024omnire , ChatSim wei2024editable , StreetGS yan2024street and test three approaches: appearance codes, standalone bilateral grids, and our multi-scale bilateral grids.
+Our focus is on modeling appearance variance across viewpoints, we build upon the open-source works DriveStudio Chen et al.
+(2024b), ChatSim Wei et al.
+(2024), StreetGS Yan et al.
+(2024a) and test three approaches: appearance codes, standalone bilateral grids, and our multi-scale bilateral grids.
 This allows for direct comparison of how each method handles photometric inconsistencies across viewpoints.
+Table 2: Comparison with different baseline approaches for scene reconstruction. Performance is reported as the average over the NuScenes Caesar et al. (2020) scenarios.
 Method Reconstruction Novel View Synthesis Geometry
 PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ CD $\downarrow$ RMSE $\downarrow$ Depth $\downarrow$
-ChatSim wei2024editable 25.10 0.805 0.252 23.27 0.725 0.270 1.557 3.509 0.106
+ChatSim Wei et al. (2024) $25.10$ $0.805$ $0.252$ $23.27$ $0.725$ $0.270$ $1.557$ $3.509$ $0.106$
 ChatSim (Ours) 27.04 0.819 0.231 24.67 0.735 0.249 1.236 3.412 0.053
-StreetGS yan2024street 25.74 0.822 0.240 23.64 0.736 0.259 1.604 3.544 0.107
+StreetGS Yan et al. (2024a) $25.74$ $0.822$ $0.240$ $23.64$ $0.736$ $0.259$ $1.604$ $3.544$ $0.107$
 StreetGS (Ours) 27.90 0.836 0.219 25.10 0.747 0.238 1.272 3.458 0.055
-Table 2: Comparison with different baseline approaches for scene reconstruction. Performance is reported as the average over the nuScenes caesar2020nuscenes scenarios.
 Implementation details.
 We evaluate our method on many autonomous driving datasets to demonstrate its robustness.
-For each dataset, We train all the methods on a single NVIDIA L40 GPU and compare the 3D reconstruction and novel view synthesis results between camera and LiDAR data.
+For each dataset, We train all the methods on a single NVIDIA L40 GPU and compare the 3D reconstruction and novel view synthesis (NVS) results between camera and LiDAR data.
 Additionally, we ablate key components of our method and quantify their impact on both appearance and geometry quality.
-Dataset Method PSNR $\uparrow$
-Full Image Human Vehicle
-Waymo sun2020scalability Baseline 26.40 23.91 21.25
-Baseline w/AC 26.44 23.81 21.34
-Baseline w/BG 21.51 22.91 19.82
-Ours 26.55 23.97 21.36
-NuScenes caesar2020nuscenes Baseline 23.74 23.55 22.80
-Baseline w/AC 23.74 23.35 22.84
-Baseline w/BG 23.60 23.63 22.79
-Ours 24.64 23.75 23.22
-Argoverse wilson2023argoverse Baseline 22.53 20.73 20.84
-Baseline w/AC 22.51 20.70 20.80
-Baseline w/BG 21.70 20.67 20.59
-Ours 22.58 20.79 20.86
-Pandaset xiao2021pandaset Baseline 27.49 26.94 21.02
-Baseline w/AC 27.51 26.98 21.13
-Baseline w/BG 27.38 27.08 20.77
-Ours 27.89 27.22 21.32
-Table 3: Comparison on Novel View Synthesis.
 Training and Dynamic Rendering.
@@
 $\mathcal{L}_{recon}=\lambda_{r}\mathcal{L}_{1}+\lambda_{s}\mathcal{L}_{\text{SSIM}}+\lambda_{d}\mathcal{L}_{d}+\lambda_{o}\mathcal{L}_{o}~~,$ (9)
-Furthermore, we introduce two additional regularization terms to enhance image fidelity: Adaptive Total Variation Regularization.
+Furthermore, we introduce two regularization terms to enhance image fidelity: Adaptive Total Variation Regularization.
 This term encourages smoothness and reduces noise while preserving image details, and Circle Regularization Loss.
-This loss applies an inverse transformation to the ground-truth images, preventing discrepancies and image quality degradation.
+This loss applies inverse transformation to the ground-truth images, preventing discrepancies and image quality degradation.
 As visually demonstrated in Fig.
-5, these two terms significantly improve image fidelity.
-Detailed explanations are provided in Appendix A1.
+5, these terms effectively improve image fidelity.
+Detailed are provided in Appendix A1.
 2) Rendering.
 For dynamic rendering using our multi-scale bilateral grid framework, we employ a specific interpolation strategy when encountering novel test images, especially those simulating dynamic ISP conditions.
-Our proposed method first conducts a temporal proximity search to identify the most relevant grids.
+We first conducts a temporal proximity search to identify the most relevant grids.
 Following this, we perform scale-specific interpolation using the two nearest grids found.
-More comprehensive details regarding this interpolation approach are available in Appendix A1.
-Figure 3: Visualizations of our method and baseline methods on Waymo sun2020scalability , NuScenes caesar2020nuscenes , Argoverse wilson2023argoverse , and PandaSet xiao2021pandaset dataset.
-Figure 4: Our proposed framework (c) outperforms appearance codes (a) and single bilateral grids (b) by simultaneously addressing optimization challenges and enhancing geometric modeling.
-This delivers significant geometric accuracy improvements (lower Chamfer Distance (CD) via multi-scale bilateral grid, reducing error and improving reconstruction) and reduced photometric inconsistency (fewer floaters) in dynamic, decoupled driving scenes.
-Yellow indicates a high lidar error, while Purple indicates a low lidar error.
+Table 3: Evaluation on Challenging Scene Subsets. Results on systematically selected extreme cases. Our method shows substantially improvements when photometric inconsistencies are severe.
+Scene Type Method Reconstruction Novel View Synthesis Geometry
+PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ CD $\downarrow$ RMSE $\downarrow$ Depth $\downarrow$
+Challenging Scenes (18 scenes) OmniRe $26.84$ $0.847$ 0.239 $24.46$ $0.758$ 0.263 $0.650$ $3.176$ $0.063$
+Ours 28.13 0.856 0.223 25.17 0.763 0.250 0.523 3.120 0.023
+Night Scenes (4 scenes) OmniRe $28.34$ $0.805$ 0.338 $26.54$ $0.753$ 0.360 $0.586$ $2.767$ $0.041$
+Ours 29.13 0.812 0.330 27.21 0.762 0.349 0.500 2.731 0.016
+Sun Flare Scenes (5 scenes) OmniRe $26.68$ $0.866$ 0.195 $24.98$ $0.796$ 0.212 $0.781$ $3.073$ $0.079$
+Ours 28.65 0.879 0.173 25.84 0.803 0.192 0.593 2.994 0.022
+Figure 3: Qualitative comparison across datasets. Our method versus baselines on Waymo, NuScenes, Argoverse, and PandaSet.
+Figure 4: Photometric consistency improves geometry.
+Our framework (c) outperforms appearance codes (a) and single bilateral grids (b) by addressing optimization challenges and enhancing geometric modeling.
+This yields lower Chamfer Distance and fewer floaters in dynamic, decoupled driving scenes.
+Yellow indicates high LiDAR error, while Purple indicates low LiDAR error.
 4.2 Quantitative Evaluation
-Reconstruction Quality.  We evaluate the reconstruction quality of our method using appearance and geometry metrics, with detailed results presented in Tab. 1.
-(1) Appearance Fidelity.
-Our approach demonstrates superior performance in appearance-based metrics (PSNR and SSIM).
+Reconstruction and Geometric Quality.
+We present a comprehensive quantitative analysis in Tab.
+1, comparing our full model against the OmniRe and its variants with appearance codes (w/AC) and a single bilateral grid (w/BG).
+Our method consistently sets a new state-of-the-art across all four datasets in both appearance and, most critically, geometric metrics.
+In terms of appearance fidelity, our model achieves the highest PSNR and SSIM scores.
+For example, on NuScenes, our method reaches a PSNR of 27.69, a significant jump from 26.37 (OmniRe) and 26.38 (OmniRe w/AC).
+This demonstrates the superior ability of the multi-scale grid to model complex photometric variations.
+However, the most striking result is the substantial improvement in geometric accuracy.
+Our method drastically reduces the Chamfer Distance, a key indicator of geometric fidelity.
+On the Waymo dataset, our model hits a CD of 0.989, outperforming the best baseline (OmniRe w/AC, 1.378) by a remarkable 28.2%.
+This trend holds across all datasets, with significant reductions in Chamfer Distance (CD), Root Mean Squared Error (RMSE), and Median Squared Depth Error (Depth).
+This result provides strong evidence for our central hypothesis: by accurately resolving photometric ambiguities, our multi-scale approach effectively eliminates geometric artifacts like "floaters" and produces a much cleaner, more accurate 3D reconstruction.
+The single bilateral grid (w/BG), despite its high expressiveness, often fails to converge and even degrades performance, highlighting the optimization challenges that our multi-scale, coarse-to-fine framework successfully overcomes.
+NVS.
 As shown in Tab.
-1, our method achieves the highest PSNR and SSIM scores for full image reconstruction across all tested datasets (Waymo, NuScenes, Argoverse, and Pandaset).
-Furthermore, when evaluating segmented dynamic regions such as humans and vehicles, our method largely surpasses the baselines, indicating its robustness in capturing both global scene appearance and fine-grained details of dynamic objects.
-For instance, on NuScenes, our PSNR for ’Vehicle’ is 27.31, compared to the best baseline score of 26.52.
-While there are isolated cases where a baseline performs comparably or slightly better (e.g., Human SSIM on Argoverse), our method generally offers the most consistent and highest quality appearance reconstruction.
-(2) Geometric Accuracy.
-In terms of geometric fidelity, our method consistently outperforms all baselines across all datasets, as evidenced by the Chamfer Distance (CD), Root Mean Square Error (RMSE), and Depth L2 Error metrics in Tab.
-1.
-For example, on the Waymo dataset, our method achieves a CD of 0.989, significantly lower than the best baseline score of 1.378.
-This superior geometric accuracy is partly attributed to our model’s improved handling of "floater" artifacts, leading to more precise LiDAR representations, as qualitatively illustrated in Fig.
-4.
-Novel View Synthesis.
-The performance of our method in novel view synthesis is detailed in Tab.
-3.
-Our approach consistently achieves the highest PSNR scores across all datasets for full images as well as for segmented human and vehicle categories.
-For example, on the NuScenes dataset, our method achieves a PSNR of 24.64 for the full image, 23.75 for humans, and 23.22 for vehicles, outperforming all baseline variations.
-This underscores the capability of our method to generate high-fidelity novel views from new camera poses.
-Comparison with State-of-the-Art Methods.
-To further assess the benefits of our architectural contributions, we integrated our core components into two recent state-of-the-art methods: ChatSim wei2024editable and StreetGS yan2024street .
-The comparative results, averaged over NuScenes scenarios, are presented in Tab.
-2.
-When our enhancements are applied (denoted as "ChatSim (Ours)" and "StreetGS (Ours)"), both methods exhibit significant improvements across all evaluated aspects.
-For instance, integrating our approach with ChatSim improves its reconstruction PSNR from 25.10 to 27.04 and its novel view synthesis PSNR from 23.27 to 24.67.
-Similarly, StreetGS sees its reconstruction PSNR boosted from 25.74 to 27.90 and its Chamfer Distance reduced from 1.604 to 1.272.
-These results highlight the generalizability and effectiveness of our proposed techniques, demonstrating their capability to enhance the performance of existing strong baseline models for comprehensive scene understanding and representation.
+1, our model’s superior reconstruction quality translates directly to improved NVS.
+Our method consistently outperforms all baselines in NVS metrics across the four datasets.
+For instance, on PandaSet, we achieve a PSNR of 27.89, surpassing the next best (OmniRe w/AC at 27.51).
+This indicates that the cleaner geometry and more consistent appearance learned by our model generalize better to unseen viewpoints, producing higher-fidelity renderings.
+Generalization to Other Methods.
+To demonstrate the broad applicability of our approach, we integrated our multi-scale bilateral grid into two other state-of-the-art methods, ChatSim Wei et al.
+(2024) and StreetGS Yan et al.
+(2024a).
+The results on NuScenes, presented in Tab.
+2, are compelling.
+Our module brings substantial gains to both frameworks.
+For StreetGS, incorporating our method boosts the reconstruction PSNR from 25.74 to 27.90 and slashes the Chamfer Distance from 1.604 to 1.272.
+These significant improvements underscore that our multi-scale grid is not just a bespoke enhancement for one architecture but a generalizable and powerful module for improving photometric and geometric consistency in Gaussian Splatting-based scene reconstruction.
+Performance on Challenging Scenarios.
+To validate robustness, we curated a subset of 18 challenging scenarios with extreme photometric inconsistencies.
+As shown in Table 3, our method’s advantages are substantially more pronounced here than on the standard datasets.
+Across this entire subset, our model achieves a +1.29 dB PSNR gain.
+This improvement is even more significant in specific difficult sub-categories; for instance, in "Sun Flare Scenes," the gain reaches +1.97 dB.
+This demonstrates our model’s strong effectiveness in resolving severe, real-world edge cases.
 4.3 Qualitative Evaluation
@@
 Fig.
-3 presents visualization results of our method on scenes from Waymo sun2020scalability , NuScenes caesar2020nuscenes , Argoverse wilson2023argoverse , and PandaSet xiao2021pandaset , demonstrating its capability to handle diverse conditions that might otherwise lead to inconsistencies.
-Figure 5: Visualizations of ablations on Circle Regularization loss and Adaptive TV loss.
-Figure 6: Visualization of Affine Transformations and Distributions.
-The first three rows display the affine transformation results and corresponding distributions when applying the first level (scale 0), the first two levels (scales 0-1), and all three levels (scales 0-1-2) of our multi-scale approach.
-The final row compares these to results from a single bilateral grid, demonstrating that the multi-scale bilateral grid achieves a coarse-to-fine, smooth image enhancement with more diverse processing capabilities.
-Affine Transformations Distributions.
-According to Fig.
-6, analyzing grayscale histograms of learned affine transformations reveals a key distinction between original single-scale and our multi-scale bilateral grids.
-For the original grid, each viewpoint’s histogram typically exhibits two distinct peaks.
-This bi-modal distribution suggests a focused, somewhat limited photometric adaptation.
-It implies that for each view, correction focuses on a few dominant transformations.
-More visualization results are shown at Fig.
-3.
-A particularly interesting observation arises examining histograms of our multi-scale grid’s coarsest level.
-Intriguingly, these histograms also demonstrate peaky distributions, varying significantly across viewpoints.
-This isn’t contradictory, but insightful: even at its coarsest level, the grid learns view-dependent base appearances.
-The coarse grid captures dominant, scene-wide photometric transformations characteristic of each viewpoint.
-Peaks in coarse-level histograms likely represent these primary, global photometric adjustments, tailored per view.
+3 and Fig.
+4 present visualization results of our method on scenes from Waymo Sun et al.
+(2020), NuScenes Caesar et al.
+(2020), Argoverse Wilson et al.
+(2023), and PandaSet Xiao et al.
+(2021), demonstrating its capability to handle diverse conditions that might otherwise lead to inconsistencies.
+Figure 5: Ablation visualizations. Effects of circle regularization and adaptive total variation losses.
+Figure 6: Affine transformation visualization.
+The first three rows show transformations and distributions for the first level, first two levels, and all three levels of our multi-scale approach.
+The final row compares against a single bilateral grid, illustrating a coarse-to-fine enhancement with richer transformation diversity.
 Configuration Grid Params (M) Training Time (h) Testing Time (FPS)
-Baseline w/AC 0.072 1.93 53
-Baseline w/BG 27.843 2.85 54 / 38†
+OmniRe w/AC 0.072 1.93 53
+OmniRe w/BG 27.843 2.85 54 / 38†
 Ours 3.969 2.10 54 / 42†
-Table 4: Computational Efficiency Analysis. †  means using bilateral grids processing during rendering.
-However, the key advantage is evident in the overlaid histogram, which aggregates distributions from all multi-scale levels and viewpoints.
-This histogram is notably flatter and more dispersed than those of the original grid, a characteristic directly resulting from hierarchical and residual refinement.
-Our framework first uses the coarse grid to establish a view-dependent base appearance.
-Subsequently, medium and fine grids learn residual transformations to correct coarse grid errors.
-Medium grids then refine regional photometric variations, while fine grids handle local, pixel-level details.
-This layered, residual approach significantly broadens the spectrum of learnable photometric transformations.
-The original grid’s two-peaked histograms signify its narrow adaptation.
-In contrast, the multi-scale version’s flatter histogram reveals its core strength: a diverse, generalizable, and robust representation of spatially-varying photometric transformations.
-This enhanced representation is crucial for consistent and high-quality renderings across diverse driving views.
+Table 4: Computational Efficiency Analysis. † denotes rendering with bilateral grid processing active.
+Analysis of Learned Affine Transformations.
+To gain deeper insight into how our model functions, we analyze the distribution of the learned affine transformations in Fig.
+6.
+The histograms for the original single-scale bilateral grid (BG) exhibit a distinct bi-modal, peaky distribution for each viewpoint.
+This suggests that the single grid learns a limited set of dominant photometric corrections for each view, struggling to capture the full spectrum of variations.
+In stark contrast, the aggregated histogram from our multi-scale grid (across all levels and viewpoints) is significantly flatter and more dispersed.
+This is a direct visualization of our framework’s core strength.
+The coarse-level grid first establishes a view-dependent baseline appearance, capturing the dominant global photometric shift (interestingly, its own histogram is also peaky, but varies across views).
+Then, the medium and fine grids learn residual transformations in a coarse-to-fine manner, correcting the errors from the preceding levels.
+This hierarchical, residual refinement process allows the model to represent a much broader and more diverse range of photometric transformations.
+The flatter histogram is empirical evidence of this enhanced representational power, which is crucial for achieving consistent, high-quality renderings across the diverse and challenging views encountered in autonomous driving.
 4.4 Ablation Studies
 As shown in Tab. 4, Tab. 5 and Fig. 5, we systematically evaluate component contributions through:
-Settings Recon.$\uparrow$ NVS.$\uparrow$ CD $\downarrow$
+[…]
   Full model 27.69 24.64 1.161
@@
 (b) Grid Size Combinations
+  Single Grid (8,8,4) 26.56 23.90 1.376
   (4,4,2)+(8,8,4) 27.49 24.56 1.963
@@
   (2,2,1)+(4,4,2) 27.33 24.42 1.213
-   (4,4,2)+(8,8,4)+(16,16,8) 27.31 24.46 2.297
 (c) Loss Functions
@@
   w/o adaptive total variance loss 27.64 24.61 1.169
-Table 5: Ablations. Numbers in Guidance Map Resolution settings indicate downsample factors from coarse level to fine level, and numbers in Grid Size Combinations settings indicate grid size.
+Table 5: Ablations. Numbers in Guidance Map Resolution indicate downsample factors from coarse to fine level, and numbers in Grid Size Combinations indicate grid size.
 Guidance Map Resolution.
-Varying guidance map resolution shows a trade-off in performance.
-Using the highest resolution (1,1,1) results in a slightly lower NVS PSNR compared to using (2,2,1) which achieves the highest NVS PSNR(24.64) in these ablations, demonstrating our patch-based approach balances efficiency and effectiveness.
+We test different downsampling factors for the guidance map at each grid level.
+The results in Tab.
+5(a) show that a moderate downsampling of (2,2,1) from coarse to fine yields the best balance, achieving the highest NVS PSNR (24.64).
+Using the full resolution (1,1,1) slightly degrades performance, likely due to overfitting to noise, validating our patch-based aggregation strategy.
 Grid Size Combinations.
-Using different combinations of grid sizes significantly impacts geometry, as indicated by CD.
-Combining smaller grids like (2,2,1)+(4,4,2) results in a lower CD (1.213) than using combinations including larger grids or the combination with the largest grid (4,4,2)+(8,8,4)+(16,16,8).
+The combination of grid sizes is crucial.
+As shown in Tab.
+5(b), removing the coarsest grid level (e.g., using only (4,4,2)+(8,8,4)) leads to a dramatic increase in CD, highlighting the importance of the coarse grid for establishing global geometric consistency.
+Crucially, a baseline with only a single fine-resolution grid (8,8,4) performs poorly, confirming that finer grids fail to converge without the stable, coarse-level initialization our method provides.
 Loss Functions.
-Ablating the circle regularization loss slightly reduces both Reconstruction PSNR (from 27.69 to 27.47) and NVS PSNR (from 24.64 to 24.52).
-Removing the adaptive total variance loss has a minimal impact on rendering result.
+Ablating our proposed regularization losses in Tab.
+5(c) shows their contribution.
+Removing the circle regularization loss slightly degrades both reconstruction and NVS performance.
+The effect of the adaptive total variance loss is more subtle in metrics but visibly improves rendering quality by reducing noise, as shown in Fig.
+5.
 Efficiency.
-As shown in Tab.
-4, while the ’Baseline w/BG’ dramatically increases parameters and training time, our method achieves a much more moderate increase in parameters and training time (to 2.10 hours) relative to the lightweight appearance codes.
-With multi-scale bilateral grid guided training, we found the rendering speed may improve due to the decreased number of unnecessary Gaussians.
+As detailed in Tab.
+4, our method strikes an effective balance between performance and computational cost.
+The single large bilateral grid (w/BG) leads to a massive parameter count (27.8M) and the longest training time (2.85h).
+In contrast, our multi-scale approach requires only 3.9M parameters and a modest 9% increase in training time (2.10h vs.
+1.93h) compared to the appearance code baseline, while delivering substantially better results.
+During inference, our method maintains high frame rates, demonstrating its practicality.
 5 Conclusion
-We propose a multi-scale bilateral grid seamlessly integrating global appearance codes and pixel-wise bilateral grids to address photometric inconsistencies in autonomous driving neural rendering.
+We introduced a novel multi-scale bilateral grid that unifies global appearance codes and pixel-wise bilateral grids into a single, hierarchical framework.
+Our extensive experiments demonstrate that this approach not only effectively models and corrects complex photometric inconsistencies in autonomous driving scenes but, more importantly, leads to a significant and crucial improvement in the geometric accuracy of the final reconstruction.
+By establishing this strong link between advanced appearance modeling and geometric fidelity, our work paves the way for more robust and reliable neural rendering systems in safety-critical, real-world applications.
 Limitations.
-While our proposed multi-scale bilateral grid demonstrates significant improvements in handling photometric inconsistencies and enhancing geometric accuracy in driving scene reconstruction, certain limitations warrant discussion.
-(1) The computational remains higher than simpler appearance code techniques; this could be a factor in highly resource-constrained deployment scenarios.
-(2) while effective for many dynamic elements, modeling extremely complex, fast-moving, or highly non-rigid objects with complete fidelity remains an ongoing challenge.
+While our method shows significant improvements, some limitations remain.
+(1) The computational overhead, though much lower than a single large bilateral grid, is still higher than simple appearance codes, which could be a consideration for resource-constrained scenarios.
+(2) Modeling extremely fast-moving or highly non-rigid objects where LiDAR and camera data misalign remains an open challenge.
+Appendix A Additional Implementation Details
+A.1 Loss Functions and Optimization
+We jointly optimize our multi-scale Gaussian scene representation by minimizing a composite loss function $\mathcal{L}_{\text{total}}$:
+$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{recon}}+\lambda_{\text{TV}}\mathcal{L}_{\text{TV}}+\lambda_{\text{circle}}\mathcal{L}_{\text{circle}}$ (10)
+where $\mathcal{L}_{\text{recon}}$ is the primary reconstruction loss, and $\mathcal{L}_{\text{TV}}$ and $\mathcal{L}_{\text{circle}}$ are regularization terms.
+A.1.1 Reconstruction Loss
+The core reconstruction loss, $\mathcal{L}_{\text{recon}}$, drives the accurate reproduction of both RGB and depth information:
+$\mathcal{L}_{\text{recon}}=\lambda_{r}\mathcal{L}_{1}+(1-\lambda_{r})\mathcal{L}_{\text{SSIM}}+\lambda_{d}\mathcal{L}_{d}+\lambda_{o}\mathcal{L}_{o}$ (11)
+Here, $\mathcal{L}_{1}$ and $\mathcal{L}_{\text{SSIM}}$ measure the image-space difference.
+$\mathcal{L}_{d}$ represents the loss between the rendered depth and the ground truth LiDAR depth.
+$\mathcal{L}_{o}$ is an opacity regularization term that encourages alignment with a non-sky mask.
+A.1.2 Adaptive Total Variation (TV) Regularization
+To encourage smoothness and reduce noise in the optimized grid representations, we incorporate a TV regularization term, $\mathcal{L}_{\text{TV}}$, at each level $l$ of the multi-scale representation:
+$\mathcal{L}_{\text{TV}}=\sum_{l}k^{(l)}\cdot\frac{1}{|\mathcal{A}^{(l)}|}\sum_{i,j,k}\sum_{\mathbb{D}\in\{x,y,z\}}\left\|\Delta_{\mathbb{D}}\mathcal{A}^{(l)}(i,j,k)\right\|_{2}^{2}$ (12)
+The adaptive weight $k^{(l)}$ for each level $l$ is proportional to the grid size, applying stronger smoothing to finer, higher-resolution grids and lighter regularization to coarser grids.
+$k^{(l)}=a\sqrt{H^{(l)}\cdot W^{(l)}\cdot D^{(l)}}+b$ (13)
+A.1.3 Circle Regularization for Photometric Consistency
+To constrain the noise introduced by photometric corrections and prevent overly aggressive alterations, we introduce a circle regularization loss, $\mathcal{L}_{\text{circle}}$:
+$\mathcal{L}_{\text{circle}}=\sum_{(u,v)\in\mathcal{S}}\left\|I^{r}(u,v)-\bar{A}^{-1}(I^{gt}(u,v))\right\|_{2}^{2}$ (14)
+where $\mathcal{S}$ denotes the set of pixel coordinates.
+This loss encourages the rendered image $I^{r}$ to be reconstructible from the ground truth image $I^{gt}$ via an inverse appearance transformation $\bar{A}^{-1}$.
+A.1.4 Coarse-to-Fine Optimization Strategy
+We employ a coarse-to-fine optimization strategy by utilizing level-dependent learning rates (e.g., $1\times 10^{-5},3\times 10^{-5},1\times 10^{-4}$ from coarse to fine).
+Coarser grids, assigned higher learning rates, rapidly learn the global scene illumination, while finer grids, with lower learning rates, hierarchically refine high-frequency photometric details.
+This staged optimization enhances stability.
+A.2 Dynamic Rendering for Real-World ISP Adaptation
+To bridge the domain gap between the fixed ISP used during training and dynamic real-world ISP pipelines encountered during inference, we employ an interpolation strategy.
+For a novel test image with timestamp $t_{\text{novel}}$, we proceed as follows:
+  1. 1.
+    Temporal Proximity Search: We identify the two training timestamps from the same camera, $t_{1}$ and $t_{2}$, that are temporally closest to $t_{\text{novel}}$ (such that $t_{1}\leq t_{\text{novel}}\leq t_{2}$).
+  2. 2.
+    Scale-Specific Grid Interpolation: We perform linear interpolation specifically on the coarse and medium-scale bilateral grids ($\mathcal{A}^{(0)}$, $\mathcal{A}^{(1)}$) derived from these two timestamps.
+The interpolated grid $\hat{\mathcal{A}}^{(l)}$ for $l\in\{0,1\}$ is formulated as:
+$\hat{\mathcal{A}}^{(l)}=\omega\mathcal{A}^{(l)}_{t_{1}}+(1-\omega)\mathcal{A}^{(l)}_{t_{2}}$ (15)
+The temporal interpolation weight, $\omega$, is determined by proximity:
+$\omega=\frac{t_{2}-t_{novel}}{t_{2}-t_{1}}$ (16)
+The fine-scale grid ($l=2$) is not used during interpolation, as it is primarily intended to capture scene-intrinsic details rather than global ISP variations.
+Appendix B Dataset Details
+This section provides dataset-specific details regarding our evaluation protocol, including sequence IDs for reproducibility.
+B.1 Waymo Open Dataset Sun et al. (2020)
+We use all five cameras and all LiDAR sensors. We conduct our experiments on the following 6 sequences, selected according to Tonderski et al. (2024); Wei et al. (2024):
+  • segment-10017090168044687777
+  • segment-10061305430875486848
+  • segment-10584247114982259878
+  • segment-15090871771939393635
+  • segment-4458730539804900192
+  • segment-5835049423600303130
+B.2 NuScenes Caesar et al. (2020)
+We utilize all six available cameras and all LiDAR sensors.
+We select the following 8 sequences: 152, 164, 171, 200, 209, 359, 529, 916, which is an extension of the dataset used by Tonderski et al.
+(2024).
+To address ego-vehicle visibility, we crop the bottom 80 pixels from the back camera images.
+B.3 PandaSet Xiao et al. (2021)
+We utilize six cameras and one LiDAR unit.
+We specifically evaluate on the following 10 challenging nighttime sequences: 063, 066, 070, 073, 074, 077, 078, 079, 088, 149.
+To mitigate ego-vehicle artifacts, we apply a bottom crop of 260 pixels to the back camera images.
+B.4 Argoverse2 Wilson et al. (2023)
+We leverage the seven ring cameras and both LiDAR sensors available in the dataset. We conduct our experiments on the following 9 sequences, in line with Tonderski et al. (2024):
+  • 05fa5048-f355-3274-b565-c0ddc547b315
+  • 0b86f508-5df9-4a46-bc59-5b9536dbde9f
+  • 185d3943-dd15-397a-8b2e-69cd86628fb7
+  • 25e5c600-36fe-3245-9cc0-40ef91620c22
+  • 27be7d34-ecb4-377b-8477-ccfd7cf4d0bc
+  • 280269f9-6111-311d-b351-ce9f63f88c81
+  • 2f2321d2-7912-3567-a789-25e46a145bda
+  • 44adf4c4-6064-362f-94d3-323ed42cfda9
+  • 5589de60-1727-3e3f-9423-33437fc5da4b
+To minimize ego-vehicle interference, we apply a bottom crop of 250 pixels to the front center, rear left, and rear right camera views.
+Appendix C Additional Quantitative Results
+This section presents a comprehensive breakdown of our model’s performance, including per-scene metrics, integration results with SOTA methods, and robustness checks on challenging subsets.
+C.1 Per-Scene Geometry and Appearance Evaluation
+We provide detailed metrics for geometry (Tab. 6 - 9) and appearance (Tab. 10 - 13).
+C.2 Integration Evaluation with SOTA Methods
+Tab. 14 and 15 provide comparisons with SOTA methods like ChatSim Wei et al. (2024) and StreetGS Yan et al. (2024a) on the NuScenes dataset.
+C.3 Performance on Challenging Scenarios
+To validate robustness, we curated subsets of challenging scenarios with extreme photometric inconsistencies (e.g., night scenes, sun glare, rain, and complex reflections).
+As shown in Tables 16, 17, and 18, our method’s performance gap widens significantly compared to all baselines (OmniRe, w/ AC, w/ BG) in these difficult conditions, demonstrating the superior robustness of our multi-scale design.
+  • Challenging Argoverse (6 scenes): 49e970c4 (Extreme Lighting, Overexposure, Lens Flare), 8184872e (Complex Shadows, Direct Sunlight), 91923e20 (Extreme Lighting, Camera Artifacts), a8a2fbc2 (Low-Angle Sun), b403f8a3 (Overpass Structure, Extreme Lighting Transitions), eaaf5ad3 (Multiple Artificial Lights, Wet Surfaces, Rainy Conditions).
+  • Challenging Waymo (6 scenes): segment-10 (Night Scene, Artificial Lighting, Lens Flare), segment-30 (Night Scene, Rainy Conditions, Specular Reflections), segment-539 (High Dynamic Range, Low-Angle Sun, Sun Glare), segment-550 (High Dynamic Range, Storefront Reflections, Hard Shadows), segment-561 (Low-Angle Sun, Lens Flare, Overexposure), segment-570 (Rainy Conditions, Low Light, Motion Blur).
+  • Challenging Pandaset (6 scenes): 19, 21, 29, 48, 52 (all Low-Angle Sun with Hard/Deep Shadows or High Dynamic Range), 63 (Night Scene, Multiple Light Sources, Glare).
+C.4 Vehicle-Specific Metrics
+To specifically evaluate performance on challenging, highly reflective surfaces such as car windows, we used semantic masks to isolate vehicles.
+Table 19 shows that our method improves not only appearance (PSNR) but also, crucially, geometry (CD) for these specific objects.
+This supports our claim that our method mitigates the negative geometric impact of such severe view-dependent effects.